# Instability and breaking of internal waves in a horizontal shear layer

### v1

@@ 5 Momentum transport
Figure 11: Top panels show time-evolving, horizontally-averaged plots of the change in the mean streamwise velocity field from the initial profile $U(y)$, for simulations $(a)$ F0.1s2.28(T), $(b)$ F1.0s5.0(T). Dotted vertical lines indicate the onset of turbulence. Arrows point to the development of a mean flow jet prior to breaking. Bottom panels show corresponding cumulative contributions to the change in the shear layer thickness $\Delta\delta_{u}$ given by the (time-integrated) terms in equation (26), averaged over the region indicated by the dashed horizontal lines in the panels above.
@@
Note that, neglecting the viscous term and using the wave polarisation relations (5), $\overline{u^{\prime}v^{\prime}}$ is initially negative (positive) when $k_{x}<0$ ($k_{x}>0$) and the right hand side is thus negative (positive) at the leading edge of the wave packet. It follows that the wave packet carries with it a negative mean flow anomaly when $k_{x}<0$ and a positive mean flow anomaly when $k_{x}>0$. This corresponds to a unidirectional Bretherton flow derived in appendix B. It is also convenient to quantify bulk changes in the momentum of the background flow in terms of the shear layer thickness $\delta_{u}$ (cf. pham2009dynamics; staquet2002transport), defined here as
@@ 5.1 Cases with $k_{x}<0$ (wave trapping)
The evolution of $\overline{u}(y)$ relative to the initial shear profile $U(y)$ is shown in figure 11(a)(a) and $(b)$ for simulations F0.1s2.28(T) and F1.0s5.0(T). The negative anomaly propagating with the wave packet is clearly seen. As the wave packet is refracted and becomes locally steep in panel $(a)$, the anomaly is focussed into a narrow jet and increases in amplitude in a manner closely resembling the self-acceleration mechanism described by sutherland2006weakly for internal waves propagating in a vertical plane. For the narrower shear layer in panel $(b)$, a thin jet also develops, though in this case the acceleration does not appear to coincide with a focussing of the wave-induced mean flow but rather emerges within the critical layer located at $y=0$ prior to the arrival of the wave packet center. These dynamics are more similar to the situation described by wang2020nonlineari with a steady small-amplitude wave forcing. The cumulative contribution of the wave-induced mean flow to the shear layer thickness is clear from panels $(c)$ and $(d)$, which reveal the competition between momentum flux from the wave that acts to decrease the shear layer width and the extraction of momentum by shear production that has the the opposite effect. In both cases, the development of turbulence after the onset of wave breaking results in a drag on the mean flow leading to a reduction in speed. This drag may be significant and in fact overcome the prior acceleration, resulting in the net increase in the shear layer thickness seen in panel $(d)$.
@@
Figure 12: Horizontally averaged profiles of the wave-induced mean streamwise flow (top row) and root mean square buoyancy perturbation field (bottom row) at successive time points indicated by the dotted to solid lines. Panels $(a)$ and $(d)$ correspond to simulation F0.1s0.76(T); $(b)$ and $(e)$ to F0.1s2.28(T); $(c)$ and $(f)$ to F1.0s5.0(T).
@@
Despite the differences in the manner of formation of the jet seen above, the similarities in its structure and the subsequent wave-breaking suggests underlying similarities in the dynamics driving instability. A key dynamical feature of the nonlinear critical layer studied by wang2020nonlineari is the local acceleration of the mean flow by the accumulation of wave pseudo-momentum (see buhler2014waves) within the critical level. Mean profiles of the root-mean-square buoyancy perturbation are shown in figure 12 alongside the adjustment to the mean flow, which show that formation of the mean flow jet defect is associated with a similarly sharp increase in the local buoyancy perturbation amplitude in all cases. For case F0.1s0.76(T) with a small amplitude, local perturbation and mean flow growth are associated with a focussing of the initial wave energy seen in panel $(a)$. On the other hand, in case F1.0s5.0(T) it is clear from panel $(c)$ that amplification occurs rapidly within the predicted critical layer location significantly prior to the arrival of the center of the wave packet. Despite these differences, the increase in both local mean flow and perturbation energy growth described by wang2020nonlineari for a steady wave forcing in unbounded stratified horizontal shear flow appears to be quite generic.
@@ 6 Energy transport
The evolution of $_{tot}, E¯\overline{E}$ and $E^{\prime}$ within the region 15<y<15-15<y<15 is plotted in figure 14 alongside the cumulative energy loss due to viscous dissipation $\int_{0}^{t}\langle\mathcal{D}+\mathcal{D}_{mean}\rangle dt$. The parameter values for the simulations shown correspond to the cases in figure 2. For simulations F0.1s2.28(T) and F1.0s5.0(T) with $k_{x}<0$ shown in panels $(a)$ and $(b)$, the total energy increases as the wave propagates into the shear layer, with positive contributions from both mean and perturbation energy. This is consistent with the expected local acceleration due to the wave-induced mean flow described above. The perturbation energy growth reaches a peak at the point where the wave starts to break, which is followed by a rapid loss of mean flow and perturbation energy due to turbulent dissipation. For weaker shear (panel $a$), the cumulative dissipation over the wave-breaking event approaches a value approximately equal to the maximum energy gain due to the wave-shear interaction, which is approximately twice the initial wave energy. However, for stronger shear (panel $b$), the cumulative dissipation is 7.5 times larger than the initial wave energy. This greatly exceeds the earlier energy growth, suggesting that the turbulent state in this case might more appropriately be thought of as being associated with a subcritical transition of the otherwise linearly stable horizontal shear flow, rather than simply balancing the total increase in energy due to the wave at the point prior to breaking. In particular, the inertial instability suggested by staquet2002transport (in which rotation provides a means for additional perturbation energy to be supplied to turbulence) does not appear to be necessary to drive significantly enhanced dissipation.
@@
A significant flux of energy out of the region 15<y<15-15<y<15 may be expected when the primary wave either passes through or is reflected, or when the wave-shear interaction and/or resulting turbulence excite further waves that radiate outwards. These effects are shown in the bottom panels of figure 14, where cumulative energy fluxes through the $y=\pm 15$ boundaries from each simulation are compared to those from an equivalent simulation with the same wave parameters but no background shear flow (cf. winters1994three). Note that in all of these cases the total flux of energy is smaller than the initial wave energy due to losses associated with (non-turbulent) viscous dissipation and dispersive effects, the former of which dominates. In simulations F0.1s2.28(T) and F1.0s5.0(T), wave trapping initially prevents the expected transmission of energy, evident from the solid blue lines in panels $(e)$ and $(f)$. It is notable however that significant generation of internal waves in F1.0s5.0(T) results in outgoing wave radiation, thereby enabling energy transmission. The decrease in the cumulative energy flux is approximately the same through the left and right boundaries suggesting wave generation is roughly symmetric. The cumulative outgoing energy flux (approximately twice the net flux through $y=15$) is almost equal to the energy of the initial wave packet and, comparing to panel $(b)$, approximately 15% of the total energy dissipated at $t=165$ (see also figure 16 and the accompanying discussion below). In simulation F6.0s0.75_kx0 (panel $g$), the wave energy transmission through the shear layer very closely matches the case with no shear, indicating that despite the significant total dissipation, there is very little energy loss from the incident wave and little to no generation of internal waves by turbulence. In simulation F6.0s0.75 (panel $h$), the reflection of the initial wave is clear, with a negative cumulative energy flux through y=15y=-15. Substantial generation of waves also occurs in this case, adding to the reflected outgoing wave energy and providing significant energy flux through $y=15$. Excluding the reflected incident wave, the total outgoing energy is approximately 20% of the total dissipation at $t=165$.
@@ 6.1 Energy partition
where diffusive fluxes have been omitted for brevity. Here the buoyancy flux $\mathcal{B}\equiv\overline{\theta^{\prime}w^{\prime}}$ is the rate of conversion of potential to kinetic energy and the shear production $\mathcal{P}_{S}\equiv-\overline{u^{\prime}v^{\prime}}({\partial\overline{u}}/{\partial y})-\overline{w^{\prime}v^{\prime}}({\partial\overline{w}}/{\partial y})$ is the rate of extraction of kinetic energy from the background flow. The cumulative evolution of these terms is shown in figure 16 alongside the cumulative kinetic and available potential energy dissipation, where terms are normalised by the total cumulative turbulent dissipation late in the simulation at $t=165$. In the wave-trapping cases shown in panels $(a)$ and $(b)$, a significant fraction of the initial energy perturbation growth through the production term is transferred to potential energy via the buoyancy flux. This fraction is noticeably smaller in panel $(b)$ when $F$ is larger. On the other hand, panels $(c)$ and $(d)$ confirm the lack of kinetic to potential energy transfer obvious from figure 15.
A significant result from figure 16 is that the cumulative dissipation rates of turbulent kinetic and potential energy depend sensitively on the energetics of the wave prior to breaking. Put simply, small-scale turbulent statistics apparently inherit properties of the larger-scale breaking wave. For example, in panels $(c)$ and $(d)$ where wave advection of momentum and hence kinetic energy dominates at the time of breaking, the cumulative dissipation at the end of the mixing event is almost entirely kinetic. On the other hand, when kinetic and potential energy is more equally partitioned at the time of breaking, as in panels $(a)$ and $(b)$, this fraction may be significantly smaller. A notable consequence is that the (cumulative) mixing efficiency $\eta(t)\equiv\int_{0}^{t}\langle\chi^{\prime}\rangle dt/\int_{0}^{t}\langle\mathcal{D}\rangle dt$ is significantly larger for these cases.
@@ 7 Discussion and conclusions
In this study, we have investigated the problem of an internal gravity wave packet propagating into a horizontal shear layer. When the wave vector forms an oblique angle with the direction of the background flow ($k_{x}\neq 0$), Doppler shifting leads to local changes in the frequency measured by a stationary observer and gradients in the flow velocity refract the wave, leading to changes in the local steepness. When $k_{x}=0$, wave propagation is unaffected by the shear layer, however significant local perturbation energy changes are possible due to advection of mean flow momentum (equivalent to the lift-up mechanism), as pointed out by bakas2009gravitya. In general, the parameter space describing the initial condition is extremely large, thus a reduced order model capturing the essential dynamical variability is desirable. Here, we constructed such a model by studying locally wave-like solutions to the ray-tracing equations under the WKBJ approximation. To provide a simple estimate from the initial conditions for the subsequent relative importance of wave advection of momentum, we derived a parameter $F$ characterising the ratio of kinetic to potential energy in the system, where large $F$ corresponds to energy perturbation growth dominated by momentum advection.
@@
The value of $F$ at a given time during the wave-shear interaction is inversely proportional to the minimum gradient Richardson number and thus, at least when the latter is positive, serves as an indicator of the susceptibility of the flow to vertical shear instability. Indeed, fully nonlinear DNS with large $F$ and no increase in wave steepness exhibited wave breaking dynamics driven by the emergence of billow-like vortices closely resembling those associated with classical Kelvin-Helmholtz instability (figures 7 and 9). Notably, this picture is also consistent with simulations of plane periodic inertia-gravity waves that, unlike in the non-rotating case, may satisfy $0<Ri_{g}<0.25$ locally and exhibit similar breaking dynamics to those observed here (lelong1998inertiaa). On the other hand, when $F$ remains small but wave refraction leads to a significant increase in the local wave steepness, wave breaking is instead predicted to occur due to an increasingly convective instability. This prediction is broadly consistent with the dynamics observed in simulations in which waves were trapped by the shear near critical levels, though questions remain regarding the role of the significant local acceleration of the mean streamwise flow in wave breaking and in particular the development of three-dimensional instabilities. The development of this flow appears to be driven by the same fundamental dynamics governing local energy growth in baroclinic critical layers under a steady wave forcing, as described by wang2020nonlineari.
@@
Figure 17: Predicted ratio of local maximum kinetic to potential energy $1+F_{\mathrm{max}}^{2}$ from ray-tracing theory versus the value obtained from DNS. Dark blue symbols correspond to simulations with initial wave amplitude $s_{0}=0.75$, whilst red symbols have $s_{0}=0.25$.
@@
The effectiveness of the linear theory for describing the local wave state prior to breaking is summarised in figure 17. In general, there is a strong positive correlation between the DNS data and the predicted local maximum ratio of kinetic to potential energy density. The agreement, though far from exact, is quite satisfactory given the assumptions of linearity and scale separation $k_{y}h\ll 1$ are generally invalid. At the very least, the use of the theory for describing qualitatively the essential dynamical variability in the wave-mean flow interaction is justified. Unsurprisingly, better agreement is found for simulations with smaller initial wave amplitude. The significant under-prediction for larger values of $F_{\mathrm{max}}$ ($k_{x}>0$) in cases where $s_{0}=0.75$ is a product of the nonlinear interaction between incident and reflected waves leading to a local increase in the wave amplitude and corresponding overturning. It worth noting here that the surprising utility of the ray-tracing theory in capturing essential dynamics of wave breaking was also noted for the problem of internal waves in a vertical shear by howland2021shear.
@@
In almost all of the DNS, energy perturbation growth eventually led to wave-breaking and turbulent dissipation. As summarised in figure 14, the dynamics of turbulent breakdown and the resulting energy exchanges between the mean flow, incident wave and turbulence depend sensitively on the route to turbulence, and hence on $F$. In both wave trapping and wave reflection scenarios, a large fraction of the total dissipation takes place during a transient burst of turbulence, much like the behaviour observed for wave breaking in a vertical shear flow (winters1994three; howland2021shear). In notable contrast however, for the horizontal shear flow studied here, larger values of $F$ are associated with a total dissipation that may be several times larger than the energy of the primary wave, thus requiring a significant input of energy to turbulence from the background flow, whose linear stability is enforced here. This may be compared with winters1994three for example, who found that the total dissipation is only around 20-25% of the initial wave energy in a vertical shear with minimum gradient Richardson number equal to $0.5$.
@@
In some cases, the wave-shear interaction led to the generation of internal waves whose total energy was seen to be up to 25% of the total energy dissipated by turbulence and more than twice that contained in the initial wave packet (figure 16). We note that even in the absence of turbulent breaking, strong generation of internal waves for $F=O(1)$ was predicted by bakas2009gravityb, who found that up to 30% of the maximum wave energy within the shear layer may be radiated away. On the other hand, localised decaying initially homogeneous and isotropic turbulence can also produce significant internal wave emission in a background stratification. Using a combination of laboratory experiments and direct numerical simulations, maffioli2014evolution found the total outgoing wave energy to be 13% of the cumulative dissipation for a horizontally localised patch of turbulence. To our knowledge, the substantial emission of internal waves by turbulence has not been reported for wave breaking at a vertical critical level (this being most obviously differentiated from the often-observed phenomenon of partial reflection by the symmetry of wave emission between each side the critical level). However, for turbulence produced in an unstable vertical shear layer in the presence of a strongly stratified region below, we can infer from pham2009dynamics that internal wave emission constituted up to 22% of the total energy dissipated.
@@
Nonetheless, the close relationship between the partition of kinetic and potential energy (which is dominated by larger scale motions) and mixing efficiency is generally consistent with results obtained from strongly stratified forced systems for which turbulence production is maintained against dissipation in a statistically stationary state by continual power input at large scales (see for example maffioli2016mixing). However, the precise relationship to the mixing efficiency is sensitive to other dimensionless turbulent parameters and, importantly, the type of forcing used. Despite broad similarities in the anisotropic structure of turbulence and the shapes of turbulent kinetic and potential energy spectra, considerable differences in mixing properties are observed between forcing concentrated in vertically rotational modes with $k_{z}=0$, forcing by internal waves and forcing with a mean vertical shear (lindborg2007stratified; howland2020mixing; portwood2019asymptotic). The results of this study indicate a wide range of possible local transient wave-breaking dynamics may contribute towards the considerable mixing variability seen in forced systems, which could in theory collectively comprise many such local wave-mean flow interactions depending on the type of forcing. Further investigation is required in order to to link robustly the local properties of wave breaking studied here to their collective influence in complex multiscale systems, where the distinction between the largest scales of turbulence and the smallest scales of internal waves is unclear.

### v2

@@ 5 Momentum transport
Figure 11: Top panels show time-evolving, $x,z$-averaged plots of the change in the mean streamwise velocity field from the initial profile $U(y)$, for simulations $(a)$ F0.1s2.28(T), $(b)$ F1.0s5.0(T). Dotted vertical lines indicate the onset of turbulence. Arrows point to the development of a mean flow jet prior to breaking. Bottom panels show corresponding cumulative contributions to the change in the shear layer thickness $\Delta\delta_{u}$ given by the (time-integrated) terms in equation (26), averaged over the region indicated by the dashed horizontal lines in the panels above.
@@
Note that, neglecting the viscous term and using the wave polarisation relations (5), $\overline{u^{\prime}v^{\prime}}$ is initially negative (positive) when $k_{x}<0$ ($k_{x}>0$) and the right hand side is thus negative (positive) at the leading edge of the wave packet. It follows that the wave packet carries with it a negative mean flow anomaly when $k_{x}<0$ and a positive mean flow anomaly when $k_{x}>0$. This corresponds to a unidirectional Bretherton flow derived in appendix C. It is also convenient to quantify bulk changes in the momentum of the background flow in terms of the shear layer thickness $\delta_{u}$ (cf. Pham, Sarkar, and Brucker, 2009; Staquet and Huerre, 2002), defined here as
@@ 5.1 Cases with $k_{x}<0$ (wave trapping)
The evolution of $\overline{u}(y)$ relative to the initial shear profile $U(y)$ is shown in figure 11(a)(a) and $(b)$ for simulations F0.1s2.28(T) and F1.0s5.0(T). The negative anomaly propagating with the wave packet is clearly seen. As the wave packet is refracted and becomes locally steep in panel $(a)$, the anomaly is focussed into a narrow jet and increases in amplitude in a manner closely resembling the self-acceleration mechanism described by Sutherland (2006b) for internal waves propagating in a vertical plane. For the narrower shear layer in panel $(b)$, a thin jet also develops, though in this case the acceleration does not appear to coincide with a focussing of the wave-induced mean flow but rather emerges within the critical layer located at $y=0$ prior to the arrival of the wave packet center. These dynamics are more similar to the situation described by Wang and Balmforth (2020) with a steady small-amplitude wave forcing. The cumulative contribution of the wave-induced mean flow to the shear layer thickness is clear from panels $(c)$ and $(d)$, which reveal the competition between momentum flux from the wave that acts to decrease the shear layer width and the extraction of momentum by shear production that has the the opposite effect. In both cases, the development of turbulence after the onset of wave breaking results in a drag on the mean flow leading to a reduction in speed. This drag may be significant and in fact overcome the prior acceleration, resulting in the net increase in the shear layer thickness seen in panel $(d)$.
@@
Figure 12: Horizontally averaged profiles of the root mean square buoyancy perturbation field (top row) and root mean square wave-induced mean streamwise flow (bottom row) at successive time points indicated by the dotted to solid lines. Panels $(a)$ and $(d)$ correspond to simulation F0.1s0.76(T); $(b)$ and $(e)$ to F0.1s2.28(T); $(c)$ and $(f)$ to F1.0s5.0(T).
@@
Despite the differences in the manner of formation of the jet seen above, the similarities in its structure and the subsequent wave-breaking suggests underlying similarities in the dynamics driving instability. A key dynamical feature of the nonlinear critical layer studied by Wang and Balmforth (2020) is the local acceleration of the mean flow by the accumulation of wave pseudo-momentum (see Bühler, 2014) within the critical level. Mean profiles of the root-mean-square buoyancy perturbation are shown in figure 12 alongside the adjustment to the mean flow, which show that formation of the mean flow jet defect is associated with a similarly sharp increase in the local buoyancy perturbation amplitude in all cases. For case F0.1s0.76(T) with a small amplitude, local perturbation and mean flow growth are associated with a focussing of the initial wave energy seen in panel $(a)$. On the other hand, in case F1.0s5.0(T) it is clear from panel $(c)$ that amplification occurs rapidly within the predicted critical layer location significantly prior to the arrival of the center of the wave packet. Despite these differences, the increase in both local mean flow and perturbation energy growth described by Wang and Balmforth (2020) for a steady wave forcing in unbounded stratified horizontal shear flow appears to be quite generic.
@@ 6 Energy transport
The evolution of $_{tot}, E¯\overline{E}$ and $E^{\prime}$ within the region 15<y<15-15<y<15 is plotted in figure 14 alongside the cumulative energy loss due to viscous dissipation $\int_{0}^{t}\langle\mathcal{D}+\mathcal{D}_{mean}\rangle dt$. The parameter values for the simulations shown correspond to the cases in figure 2. For simulations F0.1s2.28(T) and F1.0s5.0(T) with $k_{x}<0$ shown in panels $(a)$ and $(b)$, the total energy increases as the wave propagates into the shear layer, with positive contributions from both mean and perturbation energy. This is consistent with the expected local acceleration due to the wave-induced mean flow described above. The perturbation energy growth reaches a peak at the point where the wave starts to break, which is followed by a rapid loss of mean flow and perturbation energy due to turbulent dissipation. For weaker shear (panel $a$), the cumulative dissipation over the wave-breaking event approaches a value approximately equal to the maximum energy gain due to the wave-shear interaction, which is approximately twice the initial wave energy. However, for stronger shear (panel $b$), the cumulative dissipation is 7.5 times larger than the initial wave energy. This greatly exceeds the earlier energy growth, suggesting that the turbulent state in this case might more appropriately be thought of as being associated with a subcritical transition of the otherwise linearly stable horizontal shear flow, rather than simply balancing the total increase in energy due to the wave at the point prior to breaking. In particular, the inertial instability suggested by Staquet and Huerre (2002) (in which rotation provides a means for additional perturbation energy to be supplied to turbulence) does not appear to be necessary to drive significantly enhanced dissipation.
@@
A significant flux of energy out of the region 15<y<15-15<y<15 may be expected when the primary wave either passes through or is reflected, or when the wave-shear interaction and/or resulting turbulence excite further waves that radiate outwards. These effects are shown in the bottom panels of figure 14. To provide reference values for wave transmission, four additional simulations were performed with the same wave parameters but no background shear flow (cf. Winters and D’Asaro, 1994), indicated by the dashed lines in panels $e)$-$h)$. Cumulative energy fluxes through the $y=\pm 15$ boundaries from each simulation where the background shear is present are compared to those from the reference no-shear cases. Note that in all cases the total flux of energy is smaller than the initial wave energy due to losses associated with (non-turbulent) viscous dissipation and dispersive effects, the former of which dominates. In simulations F0.1s2.28(T) and F1.0s5.0(T), wave trapping initially prevents the expected transmission of energy, evident from the solid blue lines in panels $(e)$ and $(f)$. It is notable however that significant generation of internal waves in F1.0s5.0(T) results in outgoing wave radiation, thereby enabling energy transmission. The decrease in the cumulative energy flux is approximately the same through the left and right boundaries suggesting wave generation is roughly symmetric. The cumulative outgoing energy flux (approximately twice the net flux through $y=15$) is almost equal to the energy of the initial wave packet and, comparing to panel $(b)$, approximately 15% of the total energy dissipated at $t=165$ (see also figure 16 and the accompanying discussion below). In simulation F6.0s0.75_kx0 (panel $g$), the wave energy transmission through the shear layer very closely matches the case with no shear, indicating that despite the significant total dissipation, there is very little energy loss from the incident wave and little to no generation of internal waves by turbulence. In simulation F6.0s0.75 (panel $h$), the reflection of the initial wave is clear, with a negative cumulative energy flux through y=15y=-15. Substantial generation of waves also occurs in this case, adding to the reflected outgoing wave energy and providing significant energy flux through $y=15$. Excluding the reflected incident wave, the total outgoing energy is approximately 20% of the total dissipation at $t=165$.
@@ 6.1 Energy partition
where diffusive fluxes have been omitted for brevity. Here the buoyancy flux $\mathcal{B}\equiv\overline{\theta^{\prime}w^{\prime}}$ is the rate of conversion of potential to kinetic energy and the shear production $\mathcal{P}_{S}\equiv-\overline{u^{\prime}v^{\prime}}({\partial\overline{u}}/{\partial y})-\overline{w^{\prime}v^{\prime}}({\partial\overline{w}}/{\partial y})$ is the rate of extraction of kinetic energy from the background flow. The cumulative evolution of these terms is shown in figure 16 alongside the cumulative kinetic and available potential energy dissipation, where terms are normalised by the total cumulative turbulent dissipation late in the simulation at $t=165$. In the wave-trapping cases shown in panels $(a)$ and $(b)$, a significant fraction of the initial energy perturbation growth through the production term is transferred to potential energy via the buoyancy flux. This fraction is slightly reduced in panel $(b)$ when $F$ is larger. On the other hand, panels $(c)$ and $(d)$ confirm the significantly smaller fraction of kinetic to potential energy transfer obvious from figure 15.
A significant result from figure 16 is that the cumulative dissipation rates of turbulent kinetic and potential energy depend sensitively on the energetics of the wave prior to breaking. Put simply, small-scale turbulent statistics apparently inherit properties of the larger-scale breaking wave. For example, in panels $(c)$ and $(d)$ where wave advection of momentum and hence kinetic energy dominates at the time of breaking, the cumulative dissipation at the end of the mixing event is dominated by $\varepsilon$. On the other hand, when kinetic and potential energy is more equally partitioned at the time of breaking, as in panels $(a)$ and $(b)$, $\chi$ and $\varepsilon$ contribute approximately equally to the total dissipation. Note that $\chi$ is often interpreted as a diapycnal mixing rate (see, e.g. Howland et al., 2021), with a corresponding (cumulative) mixing efficiency given by $\eta(t)\equiv\int_{0}^{t}\langle\chi^{\prime}\rangle dt/\int_{0}^{t}\langle\mathcal{D}\rangle dt$. An important observed feature of the horizontally sheared system is that $\eta$ varies considerably between wave breaking scenarios, asymptoting to values from $0.08$ in panel $c)$ to $0.42$ in panel $a)$.
@@ 7 Discussion and conclusions
In this study, we have investigated the problem of an internal gravity wave packet propagating into a horizontal shear layer. When the wave vector forms an oblique angle with the direction of the background flow ($k_{x}\neq 0$), Doppler shifting leads to local changes in the intrinsic frequency measured by a stationary observer and gradients in the flow velocity refract the wave, leading to changes in the local steepness. When $k_{x}=0$, wave propagation is unaffected by the shear layer, however significant local perturbation energy changes are possible due to advection of mean flow momentum (equivalent to the lift-up mechanism), as pointed out by Bakas and Farrell (2009a). In general, the parameter space describing the initial condition is extremely large, thus a reduced order model capturing the essential dynamical variability is desirable. Here, we constructed such a model by studying locally wave-like solutions to the ray-tracing equations under the WKBJ approximation. To provide a simple estimate from the initial conditions for the subsequent relative importance of wave advection of momentum, we derived the perturbation energy ratio $F$, where large $F$ corresponds to an excess of kinetic over potential energy due to wave advection of mean flow momentum.
@@
The value of $F$ at a given time during the wave-shear interaction is inversely proportional to the square root of the minimum gradient Richardson number and thus, at least when the latter is positive, serves as an indicator of the susceptibility of the flow to vertical shear instability. Indeed, fully nonlinear DNS with large $F$ and no increase in wave steepness exhibited wave breaking dynamics driven by the emergence of billow-like vortices closely resembling those associated with classical Kelvin-Helmholtz instability (figures 7 and 9). Notably, this picture is also consistent with simulations of plane periodic inertia-gravity waves that, unlike in the non-rotating case, may satisfy $0<Ri_{g}<0.25$ locally and exhibit similar breaking dynamics to those observed here (Lelong and Dunkerton, 1998a). On the other hand, when $F$ remains small but wave refraction leads to a significant increase in the local wave steepness, large overturns in the density field are expected play a dominant role in turbulent transition (convective breaking). This prediction is broadly consistent with the dynamics observed in simulations in which waves were trapped by the shear near critical levels, though questions remain regarding the role of the significant local acceleration of the mean streamwise flow in wave breaking and in particular the development of three-dimensional instabilities. The development of this flow appears to be driven by the same fundamental dynamics governing local energy growth in baroclinic critical layers under a steady wave forcing, as described by Wang and Balmforth (2020).
@@
Figure 17: Predicted ratio of local maximum kinetic to potential energy $1+F_{\mathrm{max}}^{2}$ from ray-tracing theory versus the value obtained from DNS. Dark blue symbols correspond to simulations with initial wave amplitude $s_{0}=0.75$ (where a triangle marks the special case $k_{x}=0$), whilst red symbols have $s_{0}=0.25$.
@@
The effectiveness of the linear theory for describing the local wave state prior to breaking is summarised in figure 17. In general, there is a strong positive correlation between the DNS data and the predicted local maximum ratio of kinetic to potential energy density. The agreement, though far from exact, is quite satisfactory given the assumptions of linearity and scale separation $k_{y}h\ll 1$ are generally invalid. At the very least, the use of the theory for describing qualitatively the essential dynamical variability in the wave-mean flow interaction is justified. Unsurprisingly, better agreement is found for simulations with smaller initial wave amplitude. The under-prediction for larger values of $F_{\mathrm{max}}$ when $k_{x}>0$ in cases where $s_{0}=0.75$ (compared to smaller values of $s$ and the special case $k_{x}=0$) is a product of the nonlinear interaction between incident and reflected waves leading to a local increase in the wave amplitude and corresponding overturning. It is worth noting here that the surprising utility of the ray-tracing theory in capturing essential dynamics of wave breaking was also noted for the problem of internal waves in a vertical shear by Howland et al. (2021).
It is important to note that, whilst we have shown that linear ray-tracing theory provides reasonable predictions for the dynamics of instability, energy growth may additionally result from mechanisms not captured under the WKBJ approximation. Firstly, the WKBJ theory is invalid when $k_{y}h\ll 1$ so that the assumption of scale separation between the waves and the background flow no longer holds. In particular, this occurs near turning levels where $k_{y}\to 0$. In any case, the theory is invalid even in the linear regime, however, (linear) transient growth may still occur. These dynamics are studied formally in Arratia (2011), where the formation of layers with strong vertical shear for $k_{x}\ll 1$ is also observed. The mechanism for producing these layers, referred to as stratified lift-up, comes from wave advection of mean flow momentum, which also supplies the excess vertical shear predicted in the ray-tracing theory. Secondly, near trapping levels, $k_{y}\to\infty$, so the scale separation assumption becomes increasingly valid, however, the ray-tracing theory instead becomes invalid due to growing nonlinearities as $s$ increases. Indeed, as discussed in §5.1, evidence for nonlinear adjustments to the mean flow (through wave self-acceleration or within linear critical levels) has been observed here and warrants attention in future studies.
@@
In almost all of the DNS, energy perturbation growth eventually led to wave-breaking, turbulent dissipation and mixing. As summarised in figure 14, the dynamics of turbulent breakdown and the resulting energy exchanges between the mean flow, incident wave and turbulence depend sensitively on the route to wave breaking, and hence on $F$. In both wave trapping and wave reflection scenarios, a large fraction of the total dissipation takes place during a transient burst of turbulence, much like the behaviour observed for wave breaking in a vertical shear flow (Winters and D’Asaro, 1994; Howland et al., 2021). In notable contrast however, for the horizontal shear flow studied here, larger values of $F$ are associated with a total dissipation that may be several times larger than the energy of the primary wave, thus requiring a significant input of energy to turbulence from the background flow, whose linear stability is enforced here. The corresponding horizontal mixing of momentum results in significant deceleration of the background shear flow, as discussed in §5. This may be compared with Winters and D’Asaro (1994) for example, who found that the total dissipation is only around 20-25% of the initial wave energy in a vertical shear with minimum gradient Richardson number equal to $0.5$.
@@
In some cases, the wave-shear interaction led to the generation of internal waves whose total energy was seen to be up to 25% of the total energy dissipated by turbulence and more than twice that contained in the initial wave packet (figure 16). We note that even in the absence of turbulent breaking, strong generation of internal waves for $F=O(1)$ was predicted by Bakas and Farrell (2009b), who found that up to 30% of the maximum wave energy within the shear layer may be radiated away. On the other hand, localised decaying initially homogeneous and isotropic turbulence can also produce significant internal wave emission in a background stratification. Using a combination of laboratory experiments and direct numerical simulations, Maffioli et al. (2014) found the total outgoing wave energy to be 13% of the cumulative dissipation for a horizontally localised patch of turbulence. To our knowledge, the substantial emission of internal waves by turbulence has not been reported for wave breaking at a vertical critical level (this being most obviously differentiated from the often-observed phenomenon of partial reflection by the symmetry of wave emission between each side the critical level). However, for turbulence produced in an unstable vertical shear layer in the presence of a strongly stratified region below, we can infer from Pham et al. (2009) that internal wave emission constituted up to 22% of the total energy dissipated.
@@
Nonetheless, the close relationship between the partition of kinetic and potential energy (which is dominated by larger scale motions) and mixing efficiency is generally consistent with results obtained from strongly stratified forced systems for which turbulence production is maintained against dissipation in a statistically stationary state by continual power input at large scales (see for example Maffioli et al., 2016). However, the precise relationship to the mixing efficiency is sensitive to other dimensionless turbulent parameters and, importantly, the type of forcing used. Despite broad similarities in the anisotropic structure of turbulence and the shapes of turbulent kinetic and potential energy spectra, considerable differences in mixing properties are observed between forcing concentrated in vertically rotational modes with $k_{z}=0$, forcing by internal waves and forcing with a mean vertical shear (Lindborg and Brethouwer, 2007; Howland, Taylor, and Caulfield, 2020; Portwood, de Bruyn Kops, and Caulfield, 2019). The results of this study indicate a wide range of possible local transient wave-breaking dynamics may contribute towards the considerable mixing variability seen in forced systems, which could in theory collectively comprise many such local wave-mean flow interactions depending on the type of forcing. Further investigation is required in order to to link robustly the local properties of wave breaking studied here to their collective influence in complex multiscale systems, where the distinction between the largest scales of turbulence and the smallest scales of internal waves is unclear.